OpenStudy (anonymous):

√x+4=4 I know x=12 but how?

OpenStudy (alfie):

\[\large \sqrt{x+4} = 4 \rightarrow x+4 = 16 \rightarrow x = 12\]

OpenStudy (alfie):

This would be the second step, if you're missing that :) \[(\sqrt{x+4})^2 = (4)^2\]

OpenStudy (anonymous):

thanks alfie!